Find S24 for the sequence 2,14,...12n-10
Thank you in advance

Answer by jim_thompson5910(35256)


"2" is the first term given in the sequence
1st term = a1 = 2

The nth term is 12n-10
Plug in n = 24 to find the 24th term
24th term = a24 = 12*24-10 = 278

Now we use the arithmetic sum formula
Sn = n*(a1+an)/2
S24 = 24*(a1+a24)/2
S24 = 24*(2+278)/2
S24 = 24*(280)/2
S24 = 24*140
S24 = 3360

The sum of the first 24 terms of the sequence 2,124, ..., 12n-10 is 3360